OpenStudy (anonymous):

what percent of 5 is 3?

OpenStudy (anonymous):

let it be x % of 5 .. => x/100 * 5 = 3 Solve for x ..

OpenStudy (anonymous):

\[\frac{3}{5}=\frac{60}{100}=60\% \]